I need to solve this with either the elimination or substitution method. Thank you. It takes Boat A 3 hours longer to travel 50kms than it takes Boat B to travel 60kms. Boat A travels one third the speed of Boat B. What is the speed of each boat.
Thanks.

Answer by stanbon(48558)

solve this with either the elimination or substitution method. Thank you. It takes Boat A 3 hours longer to travel 50kms than it takes Boat B to travel 60kms. Boat A travels one third the speed of Boat B. What is the speed of each boat.
----------------------
A-Boat DATA:
distance 50 km ; Rate = x kph ; time = 50/x
-----------------------
B-Boat DATA:
distance = 60 km ; rate = 3x kph ; time = 60/(3x)
----------------------------------
EQUATION:
time A - time B = 3
50/x - 60/(3x) = 3
Multiply thru by 3x to get
150 - 60 = 9x
90 = 9x
x = 10 kph (speed of A-Boat)
3x = 30 kph (speed of B-Boat)
==========================
Cheers,
Stan H.